# Permutation polynomials with a few terms over finite fields

### v2

@@ 5. Quasi-multiplicative equivalence
Notice that $F_{i}$ for $i\in\{1,2,3,4,5,6,7,8,12,13,14\}$ are permutations for $p=3$. Since $q\equiv 2\pmod{3}$ for $f(X)$ to be a permutation, therefore, there is no need to discuss the QM equivalence between $f(X)$ and $F_{i}$ for $i\in\{1,2,3,4,5,6,7,8,12,13,14\}$. Next, we suppose that $f(X)$ is QM equivalent to $F_{9}(X)=X^{3}+aX^{q+2}+bX^{2q+1}+cX^{3q}$, where $q=5^{m}$, $c=4$, $b=a+2$, $a\neq-1$, and $m$ is odd. Then there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $21\leq d\leq q^{2}-$ 2 with $\gcd(d,q^{2}-1)=1$ such that $uf(vX^{d})=F_{9}(X)$ or equivalently,
@@
$\frac{{D_{0}}^{q+1}}{{D_{3}}^{q+1}}=\frac{(\delta^{3}-\tilde{\delta}^{q})(\delta^{3q}-\tilde{\delta})}{(\delta^{3q}-\tilde{\delta}^{q})(\delta^{3}-\tilde{\delta})}=\frac{1}{16}=1.$
@@
$\frac{{D_{0}}^{q+1}}{{D_{1}}^{q+1}}=\dfrac{-(\delta^{3}-{\tilde{\delta}}^{q})(\delta^{3q}-{\tilde{\delta}})}{(\delta^{2q+1}-{\tilde{\delta}})(\delta^{q+2}-{\tilde{\delta}}^{q})}=\frac{1}{16}=1.$
@@
This give us $(\delta^{3q}-\tilde{\delta}^{q})(\delta^{3}-\tilde{\delta})=9(\delta^{2q+1}-\tilde{\delta}^{q})(\delta^{q+2}-\tilde{\delta})$. One can take $\tilde{\delta}=-\delta^{3}$, to get $4\delta^{3q+3}=9\delta^{2q+2}(\delta+\delta^{q})^{2}$, which implies $09\delta^{2q}+9\delta^{2}+5\delta^{q+1}=$ 0. In particular, when $p=5$, we obtain that $\delta^{2q}+\delta^{2}=0$. However, it is not always true as one can choose $\delta\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ such that $\delta^{4q-4}\neq 1$. This methodology also cover the study of QM equivalence of $f(X)$ with the permutation quadrinomial $F_{18}(X)$.
Now, we show that $f(X)$ is QM inequivalent to the permutation quadrinomial $F_{19}(X)$. On the contrary, suppose $f(X)$ and $F_{19}(X)=cX^{3q}+bX^{2q+1}+aX^{q+2}+X^{3}\in\mathbb{F}_{q}[X]$ are QM equivalent. Then there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $11\leq d<q^{2}-$ 1 with $\gcd(d,q^{2}-1)=1$ such that $uf(vX^{d})=cX^{3q}+bX^{2q+1}+aX^{q+2}+X^{3}.$ Therefore, the coefficients of the polynomial $uf(vX^{d})$ are from the set $\{1,a,b,c\}$. Let $uv^{3q}D_{3}=1$, then we get that the remaining coefficients of $uf(vX^{d})$ as $\frac{D_{2}}{D_{3}}v^{-q+1},\frac{D_{1}}{D_{3}}v^{-2q+2}$ and $\frac{D_{0}}{D_{3}}v^{-3q+3}$ are in $\mathbb{F}_{q}$. Hence,
@@
when $p=3$, $b=-a$, $c=a\neq-1$, and $a^{\frac{q-1}{2}}=1$. Hence, there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $21\leq d\leq q^{2}-$ 2 with $\gcd(d,q^{2}-1)=1$ such that
@@
First, suppose that $uD_{0}v^{3}=1$, $a=uD_{3}v^{3q}$, and $b=uD_{1}v^{q+2}$. Computing $b=-a$, we get $\frac{D_{1}}{D_{3}}=-v^{2q-2}$. Raising $\frac{D_{1}}{D_{3}}=-v^{2q-2}$ to the power $q+1$, we obtain
@@
We will first see that $f(X)$ is QM inequivalent to the permutation pentanomials $G_{i}$ for $1\leq i\leq 17$, $19\leq i\leq 31$ and $i=33$ of Table LABEL:Table1. For this, we will only show the QM inequivalence of $f(X)$ with the permutation pentanomial $G_{1}(X)$ For the permutation pentanomials $G_{i}$ for $2\leq i\leq 17$, $19\leq i\leq 31$ and $i=33$, a similar approach can be followed. Let us assume that $f(X)$ is QM equivalent to $G_{1}(X)=X^{5}+X^{2^{m}+4}+X^{3\cdot 2^{m}+2}+X^{4\cdot 2^{m}+1}+X^{5\cdot 2^{m}}$ when $m\not\equiv 0\pmod{4}$. Hence, there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $21\leq d\leq q^{2}-$ 2 with $\gcd(d,q^{2}-1)=1$ such that
@@
Now, suppose $f(X)$ is QM equivalent to $G_{34}(X)$. Then, there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and an integer $21\leq d\leq q^{2}-$ 2 with $\gcd(d,q^{2}-1)=1$ such that
@@
$uv^{4q}(N_{4}+\tilde{\delta})=1,\quad a=uv^{2q+2}N_{2},\quad c=uv^{3q+1}N_{3},\quad d=uv^{4}(N_{0}+\beta^{4}\tilde{\delta}),$
@@
Here, we discuss the QM equivalence only for the first family of permutation binomials introduced in Theorem 4.17 with the permutation binomials listed in Table LABEL:Table2. One can use the similar methodology to show that the second family of permutation binomials in Theorem 4.17 is QM inequivalent to permutation binomials listed in Table LABEL:Table2. Recall that, $f(X)=(\beta^{4}\delta^{4q}+\beta^{4}{\tilde{\delta}})X^{3-q}+(\delta^{4}+{\tilde{\delta}})X^{3q-1}$ is the first family of permutation binomials we obtained. Suppose that $f(X)$ is QM equivalent to the permutation binomial $H_{1}(X)$ mentioned in Table LABEL:Table2, then there must exist a positive integer $21\leq d\leq q^{2}-$ 2 with $\gcd(d,q^{2}-1)=1$ and $u,v\in\mathbb{F}_{q^{2}}^{*}$ such that $f(X)=uH_{1}(vX^{d})$. This further implies that the set of exponents of $f(X)$ and $uH_{1}(vX^{d})$ are equal in $\mathbb{Z}_{q^{2}-1}$, that is, either $d\equiv 3-q\pmod{q^{2}-1}$ or $d\equiv 3q-1\pmod{q^{2}-1}.$ WLOG, we assume $d\equiv 3-q\pmod{q^{2}-1}$ and thus $d(q+2)\equiv 3q-1\pmod{q^{2}-1}$. This leads to the equation $(q+2)(3-q)\equiv 3q-1\pmod{q^{2}-1}$. We further obtain $2(q-1)\equiv 4\pmod{q^{2}-1}$, a contradiction to the fact that $q-1\nmid 4$ for $q>2$. Hence, $f(X)$ is QM inequivalent to the permutation binomial $H_{1}(X)$. Using the similar approach, one can see that the permutation binomials $H_{i}(X)$ for $i\in\{3,4,5,6,8\}$ listed in Table LABEL:Table2 are QM inequivalent to $f(X)$.
We next consider the permutation binomial $H_{2}(X):=X^{\frac{2^{n}-1}{2^{t}-1}+1}+aX$, $n=2^{s}t$,$s\in\{1,2\}$, $t$ is odd and $a\in\omega\mathbb{F}_{2^{t}}^{*}\cup\omega^{2}\mathbb{F}_{2^{t}}^{*}$, where $\omega$ is primitive third root of unity. Notice that for $s=1$, the sets of exponents of permutation binomials $H_{1}(X)$ and $H_{2}(X)$ are exactly same. Hence, we are done for $s=1$. Let $s=2$ and $f(X)$ is QM equivalent to $H_{2}(X)$. Therefore, the set of exponents $\{3-q,3q-1\}$ and $\{d(\frac{2^{n}-1}{2^{t}-1}+1),d\}$ of $f(X)$ and $uH_{2}(vX^{d})$, respectively, are same in $\mathbb{Z}_{q^{2}-1}$, where $21\leq d\leq q^{2}-$ 2 is an integer such that $\gcd(d,q^{2}-1)=1$. If $d\equiv 3-q\pmod{q^{2}-1}$, then we get the following expression
@@
Since $q-1$ is relatively prime to $q+1$ and $3-q$, we obtain that $(q-1)\mid 2^{t}+1=q^{1/2}+1,$ which is not possible for sufficiently large $q$. A similar argument can be applied in the case when $d\equiv 3q-1\pmod{q^{2}-1}$. Finally, we show that $f(X)$ is QM inequivalent to the permutation binomial $H_{7}(X)=X^{d^{\prime}}+aX$, $n=rk$, $d^{\prime}=\frac{2^{rk}-1}{2^{k}-1},\gcd(d^{\prime}-1,2^{k}-1)=\gcd(r,2^{k}-1)=1$ and $a\not\in\mathbb{F}_{2^{k}}^{*}$. On the contrary, let $f(X)$ and $H_{7}(X)$ be QM equivalent, which leads to existence of an interger $21\leq d\leq q^{2}-$ 2 with $\gcd(d,q^{2}-1)=1$ such that $\{3-q,3q-1\}=\{dd^{\prime},d\}$ in $\mathbb{Z}_{q^{2}-1}$. WLOG, suppose that $d\equiv 3-q\pmod{q^{2}-1}$ and $dd^{\prime}\equiv 3q-1\pmod{q^{2}-1}$, which renders
@@ 6. Conclusion
$F_{16}$ $p$ is odd, $\gcd(3,p-1)=1$,$\theta_{1}(2\theta_{4}+\theta_{3}-3\theta_{1})=\theta_{4}(\theta_{3}-\theta_{4}),\theta_{1}\in\mathbb{F}_{p^{m}}^{*},\theta_{2}\in\mathbb{F}_{p^{m}}$ and $\theta_{2}^{2}-4\theta_{1}\theta_{4}$ is a square in $\mathbb{F}_{p^{m}}^{*},$ where $\theta_{1}=a_{1}a_{3}^{p^{m}}-a_{2},\theta_{2}=a_{1}a_{2}^{p^{m}}-a_{3},\theta_{3}=a_{1}^{p^{m}+1}+a_{2}^{p^{m}+1}-a_{3}^{p^{m}+1}-1,\theta_{4}=a_{1}^{p^{m}+1}-1$ [9]
@@
$H_{2}$ $X^{\frac{2^{n}-1}{2^{t}-1}+1}+aX$ $n=2^{s}t$,$s\in\{1,2\}$, $t$ is odd and $a\in\omega\mathbb{F}_{2^{t}}^{*}\cup\omega^{2}\mathbb{F}_{2^{t}}^{*}$, where $\omega$ is primitive third root of unity [6]

### v3

@@ 5. Quasi-multiplicative equivalence
Notice that $F_{i}$ for $i\in\{1,2,3,4,5,6,7,8,12,13,14\}$ are permutations for $p=3$. Since $q\equiv 2\pmod{3}$ for $f(X)$ to be a permutation, there is no need to discuss the QM equivalence between $f(X)$ and $F_{i}$ for $i\in\{1,2,3,4,5,6,7,8,12,13,14\}$. Next, we suppose that $f(X)$ is QM equivalent to $F_{9}(X)=X^{3}+aX^{q+2}+bX^{2q+1}+cX^{3q}$, where $q=5^{m}$, $c=4$, $b=a+2$, $a\neq-1$, and $m$ is odd. Then there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $1\leq d\leq q^{2}-2$ with $\gcd(d,q^{2}-1)=1$ such that $uf(vX^{d})=F_{9}(X)$ or equivalently,
@@
$\frac{{D_{0}}^{q+1}}{{D_{3}}^{q+1}}=\frac{(\delta^{3}-\tilde{\delta}^{q})(\delta^{3q}-\tilde{\delta})}{(\delta^{3q}-\tilde{\delta}^{q})(\delta^{3}-\tilde{\delta})}=1.$
@@
$\frac{{D_{0}}^{q+1}}{{D_{1}}^{q+1}}=\dfrac{-(\delta^{3}-{\tilde{\delta}}^{q})(\delta^{3q}-{\tilde{\delta}})}{(\delta^{2q+1}-{\tilde{\delta}})(\delta^{q+2}-{\tilde{\delta}}^{q})}=1.$
@@
[…]
Now, we show that $f(X)$ is QM inequivalent to the permutation quadrinomial $F_{19}(X)$. On the contrary, suppose $f(X)$ and $F_{19}(X)=cX^{3q}+bX^{2q+1}+aX^{q+2}+X^{3}\in\mathbb{F}_{q}[X]$ are QM equivalent. Then there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $1\leq d<q^{2}-1$ with $\gcd(d,q^{2}-1)=1$ such that $uf(vX^{d})=cX^{3q}+bX^{2q+1}+aX^{q+2}+X^{3}.$ Therefore, the coefficients of the polynomial $uf(vX^{d})$ are from the set $\{1,a,b,c\}$. Let $uv^{3q}D_{3}=1$, then we get that the remaining coefficients of $uf(vX^{d})$ as $\frac{D_{2}}{D_{3}}v^{-q+1},\frac{D_{1}}{D_{3}}v^{-2q+2}$ and $\frac{D_{0}}{D_{3}}v^{-3q+3}$ are in $\mathbb{F}_{q}$. Hence,
@@
when $p=3$, b=ab=-a, $c=a\neq-1$, and $a^{\frac{q-1}{2}}=1$. Hence, there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $1\leq d\leq q^{2}-2$ with $\gcd(d,q^{2}-1)=1$ such that
@@
First, suppose that $uD_{0}v^{3}=1$, $a=uD_{3}v^{3q}$, and $b=uD_{1}v^{q+2}$. Computing b=ab=-a, we get $\frac{D_{1}}{D_{3}}=-v^{2q-2}$. Raising $\frac{D_{1}}{D_{3}}=-v^{2q-2}$ to the power $q+1$, we obtain
@@
We will first see that $f(X)$ is QM inequivalent to the permutation pentanomials $G_{i}$ for $1\leq i\leq 17$, $19\leq i\leq 31$ and $i=33$ of Table LABEL:Table1. For this, we will only show the QM inequivalence of $f(X)$ with the permutation pentanomial $G_{1}(X)$ . For the permutation pentanomials $G_{i}$ for $2\leq i\leq 17$, $19\leq i\leq 31$ and $i=33$, a similar approach can be followed. Let us assume that $f(X)$ is QM equivalent to $G_{1}(X)=X^{5}+X^{2^{m}+4}+X^{3\cdot 2^{m}+2}+X^{4\cdot 2^{m}+1}+X^{5\cdot 2^{m}}$ when $m\not\equiv 0\pmod{4}$. Hence, there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and a positive integer $1\leq d\leq q^{2}-2$ with $\gcd(d,q^{2}-1)=1$ such that
@@
Now, suppose $f(X)$ is QM equivalent to $G_{34}(X)$. Then, there exist $u,v\in\mathbb{F}_{q^{2}}^{*}$ and an integer $1\leq d\leq q^{2}-2$ with $\gcd(d,q^{2}-1)=1$ such that
@@
[…]
@@
Here, we discuss the QM equivalence only for the first family of permutation binomials introduced in Theorem 4.17 with the permutation binomials listed in Table LABEL:Table2. One can use the similar methodology to show that the second family of permutation binomials in Theorem 4.17 is QM inequivalent to permutation binomials listed in Table LABEL:Table2. Recall that, $f(X)=(\beta^{4}\delta^{4q}+\beta^{4}{\tilde{\delta}})X^{3-q}+(\delta^{4}+{\tilde{\delta}})X^{3q-1}$ is the first family of permutation binomials we obtained. Suppose that $f(X)$ is QM equivalent to the permutation binomial $H_{1}(X)$ mentioned in Table LABEL:Table2, then there must exist a positive integer $1\leq d\leq q^{2}-2$ with $\gcd(d,q^{2}-1)=1$ and $u,v\in\mathbb{F}_{q^{2}}^{*}$ such that $f(X)=uH_{1}(vX^{d})$. This further implies that the set of exponents of $f(X)$ and $uH_{1}(vX^{d})$ are equal in $\mathbb{Z}_{q^{2}-1}$, that is, either $d\equiv 3-q\pmod{q^{2}-1}$ or $d\equiv 3q-1\pmod{q^{2}-1}$. WLOG, we assume $d\equiv 3-q\pmod{q^{2}-1}$ and thus $d(q+2)\equiv 3q-1\pmod{q^{2}-1}$. This leads to the equation $(q+2)(3-q)\equiv 3q-1\pmod{q^{2}-1}$. We further obtain $2(q-1)\equiv 4\pmod{q^{2}-1}$, a contradiction to the fact that $q-1\nmid 4$ for $q>2$. Hence, $f(X)$ is QM inequivalent to the permutation binomial $H_{1}(X)$. Using the similar approach, one can see that the permutation binomials $H_{i}(X)$ for $i\in\{3,4,5,6,8\}$ listed in Table LABEL:Table2 are QM inequivalent to $f(X)$.
We next consider the permutation binomial $H_{2}(X):=X^{\frac{2^{n}-1}{2^{t}-1}+1}+aX$, $n=2^{s}t$$s\in\{1,2\}$, $t$ is odd and $a\in\omega\mathbb{F}_{2^{t}}^{*}\cup\omega^{2}\mathbb{F}_{2^{t}}^{*}$, where $\omega$ is primitive third root of unity. Notice that for $s=1$, the sets of exponents of permutation binomials $H_{1}(X)$ and $H_{2}(X)$ are exactly same. Hence, we are done for $s=1$. Let $s=2$ and $f(X)$ is QM equivalent to $H_{2}(X)$. Therefore, the set of exponents $\{3-q,3q-1\}$ and $\{d(\frac{2^{n}-1}{2^{t}-1}+1),d\}$ of $f(X)$ and $uH_{2}(vX^{d})$, respectively, are same in $\mathbb{Z}_{q^{2}-1}$, where $1\leq d\leq q^{2}-2$ is an integer such that $\gcd(d,q^{2}-1)=1$. If $d\equiv 3-q\pmod{q^{2}-1}$, then we get the following expression
@@
Since q1q-1 is relatively prime to $q+1$ and 3q3-q, we obtain that $(q-1)\mid 2^{t}+1=q^{1/2}+1,$ which is not possible for sufficiently large $q$. A similar argument can be applied in the case when $d\equiv 3q-1\pmod{q^{2}-1}$. Finally, we show that $f(X)$ is QM inequivalent to the permutation binomial $H_{7}(X)=X^{d^{\prime}}+aX$, $n=rk$, $d^{\prime}=\frac{2^{rk}-1}{2^{k}-1},\gcd(d^{\prime}-1,2^{k}-1)=\gcd(r,2^{k}-1)=1$ and $a\not\in\mathbb{F}_{2^{k}}^{*}$. On the contrary, let $f(X)$ and $H_{7}(X)$ be QM equivalent, which leads to existence of an interger $1\leq d\leq q^{2}-2$ with $\gcd(d,q^{2}-1)=1$ such that $\{3-q,3q-1\}=\{dd^{\prime},d\}$ in $\mathbb{Z}_{q^{2}-1}$. WLOG, suppose that $d\equiv 3-q\pmod{q^{2}-1}$ and $dd^{\prime}\equiv 3q-1\pmod{q^{2}-1}$, which renders
@@ 6. Conclusion
Acknowledgments
We sincerely thank the editors for handling our paper, and the referees for their careful reading, valuable comments, and constructive suggestions.
@@ Acknowledgments
$F_{16}$ $p$ is odd, $\gcd(3,p-1)=1$$\theta_{1}(2\theta_{4}+\theta_{3}-3\theta_{1})=\theta_{4}(\theta_{3}-\theta_{4}),\theta_{1}\in\mathbb{F}_{p^{m}}^{*},\theta_{2}\in\mathbb{F}_{p^{m}}$ and $\theta_{2}^{2}-4\theta_{1}\theta_{4}$ is a square in $\mathbb{F}_{p^{m}}^{*},$ where $\theta_{1}=a_{1}a_{3}^{p^{m}}-a_{2},\theta_{2}=a_{1}a_{2}^{p^{m}}-a_{3},\theta_{3}=a_{1}^{p^{m}+1}+a_{2}^{p^{m}+1}-a_{3}^{p^{m}+1}-1,\theta_{4}=a_{1}^{p^{m}+1}-1$ [9]
@@
$H_{2}$ $X^{\frac{2^{n}-1}{2^{t}-1}+1}+aX$ $n=2^{s}t$$s\in\{1,2\}$, $t$ is odd and $a\in\omega\mathbb{F}_{2^{t}}^{*}\cup\omega^{2}\mathbb{F}_{2^{t}}^{*}$, where $\omega$ is primitive third root of unity [6]